# Guided simulation of conditioned chemical reaction networks

### v1

@@ Abstract.
Let $X$ be a chemical reaction process, modeled as a multi-dimensional continuous-time jump process. Assume that at given times $0<t_{1}<\cdots<t_{n}$, linear combinations $v_{i}=L_{i}X(t_{i}),\,i=1,\dots,n$ are observed for given matrices $L_{i}$. We show how to sample the process conditioned on hitting the states $v_{1},\dots,v_{n}$ by a change of measure on the law of the unconditioned process. We derive sufficient conditions for this change of measure and complement our results using numerical illustrations.   Keywords: Chemical reaction processes, Doob’s $h$-transform, exponential change of measure, guided process.   AMS subject classification: 60J27, 60J28, 60J74
@@ 1. Introduction
Chemical reaction networks are used to study a wide class of biological, physical and chemical processes that evolve over time. For instance, one can think of the transcription of genes to mRNA and then the translation to protein, the kinetics of a virus or the dynamics of chemical components reacting with each other. The forward evolution of such processes can be described in different ways: (i) a system of ordinary differential equations, see e.g. Feinberg, (2019), (ii) a system of stochastic differential equations, see e.g. Fuchs, (2013) or (iii) continuous-time Markov jump processes, as in Anderson, (2007). It is the latter option that we consider in this paper.
@@ 1.1. Chemical reaction networks
  • a Reaction set ${\mathcal{R}}$. A reaction $\ell\in{\mathcal{R}}$ is modeled as $\sum_{k}\nu_{k\ell}S_{k}\to\sum_{k}\nu_{k\ell}^{\prime}S_{k}$. We characterize $\ell$ by the change of counts of the species $\xi_{\ell}=\left(\nu_{k\ell}^{\prime}-\nu_{k\ell}\right)_{k=1}^{n}$.
Let ${\mathbb{S}}\subseteq{\mathbb{Z}}_{+}^{d}$ denote the state space of the reaction system. A vector in ${\mathbb{S}}$ is of the form $(x_{k})_{k=1}^{d}$, where $x_{k}$ denotes the species count of type $S_{k}$. We study a Markov process $X=\left(X(t)\right)_{t\geq 0}$ on ${\mathbb{S}}$ that models the evolution of species counts over time. That is, at time $t$, the $k$-th element of $X(t)$ is given by #$\#S_{k}(t)$. At a given time $t$, the time until a reaction of type $\ell$ takes place is given by $\tau_{\ell}$. When reaction ^$\hat{\ell}=\operatornamewithlimits{argmin}_{\ell\in{\mathcal{R}}}\tau_{\ell}$ occurs, the process jumps at time $t+\tau_{\hat{\ell}}$ to $X\left(t+\tau_{\hat{\ell}}\right)=X(t)+\xi_{\hat{\ell}}$. A more detailed description of the stochastic model is given in Section 2. The process $X$ evolving on the chemical reaction network is referred to as the chemical reaction process.
@@ 1.2. Statistical problem
Suppose at fixed times $0=t_{0}<t_{1}<\cdots<t_{n}$, we observe $v_{1},\dots v_{n}$, where $v_{k}=L_{k}X(t_{k})$ with $L_{k}\in{\mathbb{R}}^{m_{k}\times d}$ and $m_{k}\leq d$, $k=1,\dots n$. We will assume the rows of each $L_{k}$ to be linearly independent. Typically reaction rates are unknown and we wish to infer those from the data. Denote the parameter vector containing those rates by $\theta$. Likelihood-based inference for $\theta$ is hampered by the lack of closed-form expressions for the transition probabilities of $X$. However, if the process were observed continuously over time, the problem would be easier. Therefore, it is natural to employ a data-augmentation scheme where we iteratively sample $X$ on $[0,t_{n}]$ conditional on $v_{1},\dots,v_{n}$ and $\theta$ and then update $\theta$ conditional on $X$. In this paper, we primarily focus on the first step, sampling from $\left(X\mid L_{k}X(t_{k})=v_{k},\,k=1,\dots n\right)$. It is a key objective of this paper to show rigorously how this can be done efficiently.
@@ 1.3. Approach: conditioning by guiding
Our approach builds on earlier work in Corstanje et al., (2023) for general Markov processes in case of a single observation. Let us highlight the main points. The law of the process $X$, conditioned to be in a given states at fixed times, is obtained through Doob’s $h$-transform. That is, there is a function $h$ that depends on the transition probabilities of $X$ that induces a change of measure such that under the new measure, ${\mathbb{P}}^{h}$, the process is conditioned to hit the observed states at times of observation. Since $h$ is typically unknown, we replace it by a fully tractable function $g$ that itself induces a change of measure to a measure ${\mathbb{P}}^{g}$. Under certain conditions, ${\mathbb{P}}^{h}$ is absolutely continuous with respect to ${\mathbb{P}}^{g}$ and
$\frac{\,\mathrm{d}{\mathbb{P}}^{h}}{\,\mathrm{d}{\mathbb{P}}^{g}}(X)=\frac{g(0,x_{0})}{h(0,x_{0})}\Psi_{t_{n}}(X).$ (1.1)
Here, $\Psi$ is known in closed form and does not depend on $h$. This result implies that $h(0,x_{0})=g(0,x_{0})\mathbb{E}^{g}\Psi_{t_{n}}(X)$. As in fact $h(0,x_{0})$ is the likelihood, we obtain an unbiased estimator for the likelihood which can for example be used in a pseudo-marginal MCMC method (Andrieu & Roberts, (2009)). A large part of this paper deals with choosing the map $(t,x)\mapsto g(t,x)$ and making the above argument precise. As we show, absolute continuity in (1.1) is not obvious and requires proof.
@@ 1.4. Related literature
In this paper, we consider discrete time observations and therefore the works below are more closely related to our work. Warne et al., (2019) give an introduction to chemical reaction processes and consider estimation for discrete time partially observation with Gaussian noise focussing on Approximate Bayesian Computation. Fearnhead, (2008) and Golightly & Sherlock, (2019) are probably closest related to our approach. The common starting points of these works is a slightly informal computation that reveals how the reaction rate of the chemical reaction processes changes upon conditioning the process on a future observation (we formalise this argument in Equation (3.4)). While the reaction rate for the conditioned process is intractable, it can be approximated and this simply boils down to choosing $g$ as above. Fearnhead, (2008) approximates $g$ using Euler discretisation of the Chemical Langevin Equation (CLE) assuming the process is fully observed without error. Golightly & Sherlock, (2019) consider conditioning on a partial observation corrupted by Gaussian noise. Their choice of $g$ is based on the linear noise approximation to the CLE. This is shown to outperform the approach of Fearnhead, (2008) and earlier work in Golightly & Wilkinson, (2015). Georgoulas et al., (2017) construct an unbiased estimator for the likelihood using random truncations and computation of matrix exponentials. This in turn is used to exploit the pseudo-marginal MCMC algorithm (Andrieu & Roberts, (2009)) for parameter estimation. Building upon this work Sherlock & Golightly, (2023) introduce the minimal extended statespace algorithm and the nearly minimal extended statespace algorithm to alleviate the problem of choosing a proposal distribution for the truncation level, as required in Georgoulas et al., (2017). Alt & Koeppl, (2023) consider the same setting as we do and derive approximations to the filtering and smoothing distributions using expectation propagation.
@@ 1.5. Contribution
We extend the result in Corstanje et al., (2023) for a single complete observation to multiple partial observations in the context of chemical reaction processes. Moreover, we discuss a variation of the next reaction algorithm by Gillespie, (1976) for sampling from a class of reaction networks with unbounded time-dependent reaction intensities. The proposed methods fit within the framework of Backward Filtering Forward Guiding (van der Meulen & Schauer, (2020)), drawing strongly on techniques for exponential changes of measure as outlined in Palmowski & Rolski, (2002). This enables to construct a guided process that at any time takes into account all future conditionings.
Compared to Fearnhead, (2008) and Golightly & Sherlock, (2019), we consider the setting of multiple future conditionings (rather than one), without imposing extrinsic noise on the observations. Moreover, we derive the conditioned process and likelihood ratio in (1.1) on path space. Sufficient conditions on $g$ to guarantee absolute continuity are given in Theorem 3.7. It turns out that the choices for $g$ in Fearnhead, (2008) and Golightly & Sherlock, (2019) satisfy the assumptions of this theorem. In numerical examples we show that flexibility in choosing $g$ is particularly beneficial in case some of the components of the chemical reaction process have counts that vary monotonically over time.
@@ 1.6. Outline
We introduce stochastic chemical reaction processes in Section 2 and discuss examples that we will study. In Section 3, we describe the approach of conditioning chemical reaction processes by guiding through Doob’s $h$-transform and present conditions on $g$ such that ${\mathbb{P}}^{h}\ll{\mathbb{P}}^{g}$. In Section 4 we consider various choices for $g$. In Section 5 we present methods for simulation of conditioned chemical reaction processes together with numerical illustrations. We end with a discussion section. The appendix contains various proofs.
@@ 1.7. Frequently used notation
Throughout, we assume that we have an underlying right-continuous filtered probability space $\left(\Omega,{\mathcal{F}},\{{\mathcal{F}}_{t}\}_{t},{\mathbb{P}}\right)$ such that ${\mathcal{F}}=\bigvee_{t}{\mathcal{F}}_{t}$, $\{{\mathcal{F}}_{t}\}_{t}$ is a right-continuous filtration and all processes are adapted to the filtration. The restriction of a probability measure ${\mathbb{P}}$ or expectation ${\mathbb{E}}$ to ${\mathcal{F}}_{t}$ is denoted by ${\mathbb{P}}_{t}$ or ${\mathbb{E}}_{t}$, respectively. For a Markov process $X$, with law ${\mathbb{P}}$, we denote the measure induced by a Doob $h$-transform using a function $g$ by ${\mathbb{P}}^{g}$. Expectations with respect to ${\mathbb{P}}^{g}$ is denoted by ${\mathbb{E}}^{g}$. The generator of $X$ is denoted by ${\mathcal{L}}$ and we refer to ${\mathcal{A}}=\frac{\partial}{\partial t}+{\mathcal{L}}$ as the space-time generator. Under ${\mathbb{P}}^{g}$, these are ${\mathcal{L}}^{g}$ and ${\mathcal{A}}^{g}$. The notation is summarized in the table below.
Intuition Law Expectation Space-time generator
original Markov process ${\mathbb{P}}_{t}$ ${\mathbb{E}}_{t}$ ${\mathcal{A}}$
transformed process induced by $g$ ${\mathbb{P}}^{g}_{t}$ ${\mathbb{E}}_{t}^{g}$ ${\mathcal{A}}^{g}$
Furthermore, for a stochastic process $X$, we use the notation $X^{t}=\{X(s)\colon s\leq t\}$. Given $L\in{\mathbb{R}}^{m\times d}$ with $m\leq d$ and $v\in{\mathbb{R}}^{m}$, we denote the inverse image of $v$ under $L$ by $L^{-1}v=\{x\in{\mathbb{R}}^{d}:Lx=v\}$. For functions $f_{1}(t,x)$ and $f_{2}(t,x)$ of time and space, we say that $f_{1}\propto f_{2}$ if there exists a differentiable function $\kappa$ of time such that for all $t,x$, $f_{1}(t,x)=\kappa(t)f_{2}(t,x)$. Derivatives with respect to a variable representing $t$, say $\partial/\partial t$ are denoted by $\partial_{t}$. We denote by
$A_{n}=\{L_{k}X(t_{k})=v_{k},\,k=1,\dots,n\}.$
@@ 2. Chemical reaction processes
We construct a stochastic process to model the dynamics of the chemical reaction network described in Section 1.1 following Chapter 1 of Anderson & Kurtz, (2015). Let $X$ be a Markov process on ${\mathbb{S}}$ such that the $i$’th component of $X(t)$, $X_{i}(t)$, represents the concentration of species $S_{i}$ at time $t$. A reaction $\ell$ is represented through a difference vector $\xi_{\ell}\in{\mathbb{S}}$ and an intensity $\lambda_{\ell}\colon[0,\infty)\times{\mathbb{S}}\to[0,\infty)$. That is, $\lambda_{\ell}$ and $\xi_{\ell}$ are such that
${\mathbb{P}}\left(X({t+\Delta})-X(t)=\xi_{\ell}\mid{\mathcal{F}}_{t}\right)=\lambda_{\ell}(t,X(t))\Delta+o(\Delta),\qquad\Delta\downarrow 0.$ (2.1)
The jump probabilities specified in (2.1) correspond to a jump process with jumps $(\xi_{\ell})_{\ell\in{\mathcal{R}}}$ and exponential jump rate functions $(\lambda_{\ell})_{\ell\in{\mathcal{R}}}$. The space-time process $(t,X_{t})$ has infinitesimal generator
${\mathcal{A}}f(t,x)=\partial_{t}f(t,x)+{\mathcal{L}}_{t}f(t,x),$
where
${\mathcal{L}}_{t}f(t,x)=\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(t,x)\left[f(t,x+\xi_{\ell})-f(t,x)\right],\qquad f\in{\mathcal{D}}({\mathcal{L}}_{t}),\>x\in{\mathbb{S}},.$ (2.2)
where, the domain ${\mathcal{D}}({\mathcal{L}}_{t})$ of ${\mathcal{L}}_{t}$ consists of all bounded functions $f\colon{\mathbb{S}}\to{\mathbb{R}}$. Throughout, we impose the following assumptions on the network.
@@
    For all $t\geq 0$ and $x\in{\mathbb{S}}$, $\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(t,x)<\infty$.
  4. (2.1d)
    For $f$ with finite support in ${\mathbb{S}}$, $t\geq 0$ and ${\mathcal{L}}_{t}$ defined in (2.2), $\lim_{\left\lvert x\right\rvert\to\infty}{\mathcal{L}}_{t}f(x)=0$.
In particular, $x\mapsto\lambda_{\ell}(t,x)$ need not be linear. For a stochastic process $X(t)$, let $T_{K}=\inf\{t:|X(t)|\geq K\}$ and $T_{\infty}=\lim_{K\to\infty}T_{K}$.
Theorem 2.2 (Theorem 1.22 of Anderson & Kurtz, (2015)).
Suppose 2.1 is satisfied and let $X$ be a jump process with jumps $(\xi_{\ell})_{\ell\in{\mathcal{R}}}$ and exponential jump rate functions $(\lambda_{\ell})_{\ell\in{\mathcal{R}}}$ satisfying
@@
where $(Y_{\ell})_{\ell\in{\mathcal{R}}}$ are independent, unit rate Poisson processes. Set $X(t)=\infty$ if $t\geq T_{\infty}$. Then the solution to (2.3) is the unique minimal solution of the martingale problem for (2.2). That is, there exists a filtration $\{{\mathcal{F}}_{t}\}$ such that for each $f$ with finite support
$f(t,X(t))-\int_{0}^{t}{\mathcal{A}}f(s,X(s))\,\mathrm{d}s$
is a $\{{\mathcal{F}}_{t}\}$-martingale.

### v2

@@ Abstract.
Let $X$ be a chemical reaction process, modeled as a multi-dimensional continuous-time jump process. Assume that at given times $0<t_{1}<\cdots<t_{n}$, linear combinations $v_{i}=L_{i}X(t_{i}),\,i=1,\dots,n$ are observed for given matrices $L_{i}$. We show how the process that is conditioned on hitting the states $v_{1},\dots,v_{n}$ is obtained by a change of measure on the law of the unconditioned process. This results in an algorithm for obtaining weighted samples from the conditioned process. Our results are illustrated by numerical simulations.
Keywords: Chemical reaction processes, Doob’s $h$-transform, exponential change of measure, guided process.
AMS subject classification: 60J27, 60J28, 60J74
@@ 1. Introduction
Chemical reaction networks are used to study a wide class of biological, physical and chemical processes that evolve over time. For instance, one can think of the transcription of genes to mRNA and then the translation to protein, the kinetics of a virus or the dynamics of chemical components reacting with each other. The forward evolution of such processes can be described in different ways: (i) a system of ordinary differential equations, see e.g. Feinberg, (2019), (ii) a system of stochastic differential equations, see e.g. Fuchs, (2013) or (iii) continuous-time Markov jump processes, as in Anderson, (2007). It is the third option that we consider in this paper.
@@ 1.1. Chemical reaction networks
  • a Reaction set ${\mathcal{R}}$. A reaction $\ell\in{\mathcal{R}}$ is modeled as $\sum_{k}\nu_{k\ell}S_{k}\to\sum_{k}\nu_{k\ell}^{\prime}S_{k}$. We characterize $\ell$ by the change of counts of the species $\xi_{\ell}=\left(\nu_{k\ell}^{\prime}-\nu_{k\ell}\right)_{k=1}^{d}$.
Let ${\mathbb{S}}\subseteq{\mathbb{Z}}_{\geq 0}^{d}$ denote the state space of the reaction system. A vector in ${\mathbb{S}}$ is of the form $(x_{k})_{k=1}^{d}$, where $x_{k}$ denotes the species count of type $S_{k}$. We study a Markov process $X=\left(X(t)\right)_{t\geq 0}$ on ${\mathbb{S}}$ that models the evolution of species counts over time. That is, at time $t$, the $k$-th element of $X(t)$ is given by #$\#S_{k}(t)$. At a given time $t$, the time until a reaction of type $\ell$ takes place is given by $\tau_{\ell}$. When reaction ^$\hat{\ell}=\operatornamewithlimits{argmin}_{\ell\in{\mathcal{R}}}\tau_{\ell}$ occurs, the process jumps at time $t+\tau_{\hat{\ell}}$ to $X\left(t+\tau_{\hat{\ell}}\right)=X(t)+\xi_{\hat{\ell}}$. A more detailed description of the stochastic model is given in Section 2. The process $X$ evolving on the chemical reaction network is referred to as the chemical reaction process.
@@ 1.2. Statistical problem
Suppose at fixed times $0=t_{0}<t_{1}<\cdots<t_{n}$, we observe $v_{1},\dots v_{n}$, where $v_{k}=L_{k}X(t_{k})$ with $L_{k}\in{\mathbb{R}}^{m_{k}\times d}$ and $m_{k}\leq d$, $k=1,\dots n$. Not assuming $L$ to be the identity matrix is for example important in applications where the measuring device cannot distinguish two or more species, so that only sums of their counts are observed. We will assume the rows of each $L_{k}$ to be linearly independent. Typically reaction rates are unknown and we wish to infer those from the data. Denote the parameter vector containing those rates by $\theta$. Likelihood-based inference for $\theta$ is hampered by the lack of closed-form expressions for the transition probabilities of $X$. However, if the process were observed continuously over time, the problem would be easier. Therefore, it is natural to employ a data-augmentation scheme where we iteratively sample $X$ on $[0,t_{n}]$ conditional on $v_{1},\dots,v_{n}$ and $\theta$ and then update $\theta$ conditional on $X$. In this paper, we focus on the first step, sampling from $\left(X\mid L_{k}X(t_{k})=v_{k},\,k=1,\dots n\right)$. It is a key objective of this paper to show rigorously how this can be done efficiently. Note that a simple rejection sampling scheme where we discard all paths contradicting the observations is valid but very inefficient in most settings.
@@ 1.3. Approach: conditioning by guiding
Our approach builds on earlier work in Corstanje et al., (2023) for general Markov processes in case of a single observation. Let us highlight the main points. The law of the process $X$, conditioned to be in a given state at fixed times, is obtained through Doob’s $h$-transform. That is, there is a function $h$ that depends on the transition probabilities of $X$ that induces a change of measure such that under a new measure, ${\mathbb{P}}^{h}$, the process is conditioned to hit the observed states at times of observation. Since $h$ is typically unknown, we replace it by a fully tractable function $g$ that itself induces a change of measure to a measure ${\mathbb{P}}^{g}$. Under certain conditions, ${\mathbb{P}}^{h}$ is absolutely continuous with respect to ${\mathbb{P}}^{g}$ and
$\frac{\,\mathrm{d}{\mathbb{P}}^{h}}{\,\mathrm{d}{\mathbb{P}}^{g}}(X)=\frac{1}{h(0,x_{0})}F(X).$ (1.1)
Here, $F$ is known in closed form, depends on $g$ but does not depend on $h$. Weighted samples of the conditioned process can therefore be obtained by sampling under ${\mathbb{P}}^{g}$. The argument for making the above precise is not too hard in case $g$ is bounded and bounded away from zero. However, some natural choices we discuss and have been proposed in the literature require a more delicate argument.
@@ 1.4. Related literature
In this paper, we consider discrete-time observations and therefore the works below are more closely related to our work. Warne et al., (2019) give an introduction to chemical reaction processes and consider estimation for discrete-time partial observations with Gaussian noise focussing on Approximate Bayesian Computation. Fearnhead, (2008) and Golightly & Sherlock, (2019) are probably closest related to our approach. The common starting points of these works is a slightly informal computation that reveals how the reaction rate of the chemical reaction processes changes upon conditioning the process on a future observation (we present this argument at the start of Section 3). While the reaction rate for the conditioned process is intractable, it can be approximated and this simply boils down to choosing $g$ as above. Fearnhead, (2008) approximates $g$ using Euler discretisation of the Chemical Langevin Equation (CLE) assuming the process is fully observed without error. Golightly & Sherlock, (2019) consider conditioning on a partial observation corrupted by Gaussian noise. Their choice of $g$ is based on the linear noise approximation to the CLE. This is shown to outperform the approach of Fearnhead, (2008) and earlier work in Golightly & Wilkinson, (2015).
Georgoulas et al., (2017) construct an unbiased estimator for the likelihood using random truncations and computation of matrix exponentials. This in turn is used to exploit the pseudo-marginal MCMC algorithm (Andrieu & Roberts, (2009)) for parameter estimation. Building upon this work Sherlock & Golightly, (2023) introduce the minimal extended statespace algorithm and the nearly minimal extended statespace algorithm to alleviate the problem of choosing a proposal distribution for the truncation level, as required in Georgoulas et al., (2017).
Alt & Koeppl, (2023) consider the same setting as we do and derive approximations to the filtering and smoothing distributions using expectation propagation.
@@ 1.5. Contribution
We provide sufficient conditions on $g$ such that (1.1) holds true. We extend the result in Corstanje et al., (2023) for a single complete observation to multiple partial observations in the context of chemical reaction processes. Moreover, we discuss a variation of the next reaction algorithm by Gillespie, (1976) for sampling from a class of reaction networks with unbounded time-dependent reaction intensities.
The proposed methods fit within the framework of Backward Filtering Forward Guiding (van der Meulen & Schauer, (2020)), drawing strongly on techniques for exponential changes of measure as outlined in Palmowski & Rolski, (2002). This enables to construct a guided process that at any time takes into account all future conditionings.
Compared to Fearnhead, (2008) and Golightly & Sherlock, (2019), we consider the setting of multiple future conditionings (rather than one), without imposing extrinsic noise on the observations. Moreover, we derive the conditioned process and likelihood ratio in (1.1) on path space. Sufficient conditions on $g$ to guarantee absolute continuity are given in Theorem 3.4. It turns out that the choices for $g$ in Fearnhead, (2008) and Golightly & Sherlock, (2019) satisfy the assumptions of this theorem. In numerical examples we show that flexibility in choosing $g$ is particularly beneficial in cases where some of the components of the chemical reaction process have counts that vary monotonically over time.
@@ 1.6. Outline
We introduce stochastic chemical reaction processes in Section 2 and discuss examples that we will study. In Section 3, we describe the approach of conditioning chemical reaction processes by guiding and present conditions on $g$ such that ${\mathbb{P}}^{h}$ is absolutely continuous with respect to ${\mathbb{P}}^{g}$. In Section 4 we consider various choices for $g$. Conditions for equivalence of ${\mathbb{P}}^{h}$ and ${\mathbb{P}}^{g}$ are discussed in Section 5. In Section 6 we present methods for simulation of conditioned chemical reaction processes together with numerical illustrations. We end with a discussion section. The appendix contains various proofs.
@@ 1.7. Frequently used notation
Throughout, we assume that we have an underlying probability space $\left(\Omega,{\mathcal{F}},{\mathbb{P}}\right)$. For a stochastic process $X$, we use the notation $X^{t}=\{X(s)\colon s\leq t\}$. Given $L\in{\mathbb{R}}^{m\times d}$ with $m\leq d$ and $v\in{\mathbb{R}}^{m}$, we denote the inverse image of $v$ under $L$ by $L^{-1}v=\{x\in{\mathbb{R}}^{d}:Lx=v\}$. For functions $f_{1}(t,x)$ and $f_{2}(t,x)$ of time and space, we say that $f_{1}\propto f_{2}$ if there exists a differentiable function $\kappa$ of time such that for all $t,x$, $f_{1}(t,x)=\kappa(t)f_{2}(t,x)$. Derivatives with respect to a variable representing $t$, say $\partial/\partial t$ are denoted by $\partial_{t}$. We denote by
$A_{n}=\{L_{k}X(t_{k})=v_{k},\,k=1,\dots,n\},$ (1.2)
@@ 2. Chemical reaction processes
We construct a stochastic process to model the dynamics of the chemical reaction network described in Section 1.1 following Chapter 1 of Anderson & Kurtz, (2015). Let $X$ be a Markov process on ${\mathbb{S}}$ such that the $i$-th component of $X(t)$, $X_{i}(t)$, represents the frequency of species $S_{i}$ at time $t$. A reaction $\ell$ is represented through a difference vector $\xi_{\ell}\in{\mathbb{S}}$ and an intensity $\lambda_{\ell}\colon[0,\infty)\times{\mathbb{S}}\to[0,\infty)$. We assume
${\mathbb{P}}\left(X({t+\Delta})-X(t)=\xi_{\ell}\mid{\mathcal{F}}_{t}^{X}\right)=\lambda_{\ell}(t,X(t))\Delta+o(\Delta),\qquad\Delta\downarrow 0,$ (2.1)
where ${\mathcal{F}}_{t}^{X}=\sigma(X_{s},\,s\leq t)$. The jump probabilities specified in (2.1) correspond to a process with jumps $(\xi_{\ell})_{\ell\in{\mathcal{R}}}$ and jump rate functions $(\lambda_{\ell})_{\ell\in{\mathcal{R}}}$. Throughout, we impose the following assumptions on the network.
@@
    For all $t\geq 0$ and $x\in{\mathbb{S}}$, $\int_{0}^{t}\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(s,x)\,\mathrm{d}s<\infty$.
For a stochastic process $X(t)$, let $T_{K}=\inf\{t:|X(t)|\geq K\}$ and $T_{\infty}=\lim_{K\to\infty}T_{K}$. Let $X$ be the jump process with jumps $(\xi_{\ell})_{\ell\in{\mathcal{R}}}$ and jump rate functions $(\lambda_{\ell})_{\ell\in{\mathcal{R}}}$ satisfying
@@
where $(Y_{\ell})_{\ell\in{\mathcal{R}}}$ are independent, unit rate Poisson processes. We assume the process to be non-explosive: $\mathbb{P}(T_{\infty}<\infty)=0$. For $f\colon[0,T]\times{\mathbb{S}}\to{\mathbb{R}}$ with $f(t,\cdot)$ finitely supported for every $t$ define
$({\mathcal{L}}f)(t,x)=\sum_{\ell\in{\mathcal{R}}}\lambda_{\ell}(t,x)\left[f(t,x+\xi_{\ell})-f(t,x)\right],$ (2.3)
By Theorem 1.22 in Anderson & Kurtz, (2015), there exists a filtration $({\mathcal{F}}_{t})_{t}$ such that for all such $f$
$D^{f}(t):=f(t,X(t))-\int_{0}^{t}(\partial_{t}+{\mathcal{L}})f(s,X(s))\,\mathrm{d}s$
is an $\{{\mathcal{F}}_{t}\}$-martingale. That is, $X$ is the unique solution to the martingale problem for ${\mathcal{A}}:=\partial_{t}+{\mathcal{L}}$.